# Modular transformations of on-shell actions of (root-)$\overline{\text{T}}$ deformed holographic CFTs

### v2

@@ 4.2 Thermodynamics
After the $\overline{\text{T}}$ deformation, the entropy of the black hole does not get deformed, thus it is equal to101010which is also equal to the black hole entropy of the undeformed theory.
@@
which not only depends on the modular parameter but also depends on the size of the system. For example, the deformation is insignificant when the torus is infinite long. This is because that the effective deformation parameter is the dimensionless quantity $\rho_{c}/|\bm{\omega}_{\mu}|^{2}$. As expected, the deformed on-shell action still equals to the free energy
@@
even though both sides of this equation are deformed in a non-trivial way. To examine the first law of thermodynamics we treat $M_{\mu}$, $J_{\mu}$ and $S_{BH}^{[\mu]}$ as functions of $\bm{\beta}_{\mu}$ and $\bm{\bar{\beta}}_{\mu}$ and do the variation with respect to them. It turns out the first law is in the same form of the undeformed one
@@ 4.3 Hawking-Page transition and Modular transformation
By equating them, we find that Hawking-Page phase transition point remains unchanged at $|\bm{\tau}_{\mu}|=1$. Clearly, the deformed on-shell actions $I_{E,\text{global}}^{[\mu]}$ and $I_{E,\text{BTZ}}^{[\mu]}$ are interrelated through the following modular transformation
@@
This transformation aligns with the modular property exhibited by the $\overline{\text{T}}$ deformed torus partition function, as proved in [29]. To further support our argument regarding modular transformation, let us proceed to compute the deformed on-shell action for other $SL(2,\mathbb{Z})$ black holes. It is noteworthy that since the $\overline{\text{T}}$ deformation alters the temperature, the modular transformation does not commute with the $\overline{\text{T}}$ deformation. In other words,
@@
and making the following transformation
@@
we end up with the deformed $SL(2,\mathbb{Z})$ black hole metric which is in the form (2) with
@@
$\bm{\hat{\omega}}_{\mu}=\bm{\omega}_{\mu}(a_{m}\bm{\tau}_{\mu}+b_{m}),\quad\bm{\hat{\beta}}_{\mu}=\bm{\omega}_{\mu}(c_{m}\bm{\tau}_{\mu}+d_{m})$ (4.37)
as expected for the $SL(2,{\mathbb{Z}})$ black hole. Since the coordinate transformation (4.35) will not change the on-shell action, we conclude that
@@
which is exactly the modular transformation found in [29]. From the holographic perspective, the modular transformation can be readily understood through the transformation given in equation (4.35). The deformation is irrelevant, and as a result, the deformation parameter $\mu$ has a dimension and scales with the radial coordinate $z$ under a Weyl transformation.
@@
as the modular transformation simply maps one saddle to another. However, this semi-classical partition function is not holomorphic factorized because $\tau_{\mu}$ and $\bar{\tau}_{\mu}$ transform collectively. To further illustrate that the modular property persists even when the holomorphic and anti-holomorphic sectors are asymmetric, we will explore the $\overline{\text{T}}$ deformation of topologically massive gravity, which is dual to a conformal field theory with distinct holomorphic and anti-holomorphic central charges, i.e. $c_{L}\neq c_{R}$[94, 95]. The AdS/CFT correspondence for TMG is elegantly established in [108]. For our purposes of computing the deformed on-shell action, we solely require the expression of the holographic energy-stress tensor to define the $\overline{\text{T}}$ deformation, as we shall demonstrate subsequently.
@@ 5 Topologically massive gravity theory
We consider the Euclidean version of the TMG theory whose action is given by
@@
where $\mathcal{L}_{CS}$ is the gravitational Chern-Simons term. Notably, we assume the parameter $\alpha$ to be real, ensuring that the deformed metric remains real [^12] For the Euclidean TMG, derived through a Wick rotation from the Lorentzian theory, the parameter $\alpha$ is purely imaginary because $\mathcal{L}_{CS}$ has odd parity.. The equation of motion for this theory reads
@@
Consequently, the Einstein solutions (2.11) also satisfy the equations of TMG. However, the holographic stress tensor is modified to [95, 103]
@@
where $\epsilon_{ij}$ is the two-dimensional Levi-Civita tensor. The actions of global AdS and BTZ black holes in TMG are given in [109, 111]:
@@
where the involved parameter $k$ and $\bar{k}$ is determined by the relation
@@ 5.1 $\overline{\text{T}}$ deformation
therefore the flow equation is still solved by (2.4) and (2.5). Substituting the stress tensor (5.5) into (2.4) leads to
$\gamma_{\alpha\beta}^{[\mu]}dx^{\alpha}dx^{\beta}=(dv+\frac{\mu\bar{k}}{4\pi}\bar{L}_{0}d\bar{v})(d\bar{v}+\frac{\mu k}{4\pi}L_{0}dv)\equiv dVd\bar{V},$ (5.9)
@@
Notice that (5.9) can not be interpreted as an induced metric on a cut-off surface, hence, the cut-off (or glue-on) proposal for $\overline{\text{T}}$ deformation is no longer valid. The deformed on-shell action consists of two parts (2.9) which are convenient to express in terms of quantities in the seed theory:
@@
where $A_{0}$ is the area of the seed torus. In the end, we will use the dictionary (5.10) to rewrite it in terms of the quantities in the deformed theory. Until now, our discussion is very general. It is enough to demonstrate the modular transformation between deformed global AdS and deformed BTZ and other case can be generated trivially as we did in section 4.3.
@@
It has been demonstrated in Section 4.3, the change of the deformation parameter under the modular transformation can be understood from the fact that the $\overline{\text{T}}$ deformation is irrelevant. We proceed to investigate the modular properties of the root-$\overline{\text{T}}$ deformation, which is classically marginal. It is worth noting that the root-$\overline{\text{T}}$ deformation is both compatible with the $\overline{\text{T}}$ deformation and also preserves classical integrability [83]. Although the quantum aspects of root-$\overline{\text{T}}$ deformation are less understood [^14] See [84] for some recent progress., a holographic description has been proposed in [85], which bears similarity to the mixed-boundary-condition description of the $\overline{\text{T}}$ deformation. In particular, the spectrum of the deformed theory is accurately reproduced. The authors of [85] also suggest that the root-$\overline{\text{T}}$ deformed CFT might possess certain modular properties. In particular, in [79] it is shown that S-modular transformation is preserved. We provide some evidences by evaluating the deformed on-shell action, as we did for the $\overline{\text{T}}$ deformation.
@@ 6 Root-$\overline{\text{T}}$ deformation
The operator $\mathcal{R}$ is a multi-trace operator, its holographic dual also corresponds to a deformation of the boundary metric. Because the operator is marginal, it is no need to add a boundary term to the on-shell action like the one in (2.9). The solution of the flow equation is not unique but it is argued in [85], the proper root-$\overline{\text{T}}$ deformed boundary metric and stress tensor are given by
@@
where we have used the fact that for a classically marginal deformation, the trace of the stress tensor does not flow. Again we can introduce the flat coordinates on the deformed boundary metric
$\gamma_{\alpha\beta}^{(\mu)}dx^{\alpha}dx^{\beta}=dVd\bar{V}$ (6.6)
@@
$dV=\cosh(\frac{\mu}{2})dv+\sqrt{\frac{\bar{L}_{0}}{L_{0}}}\sinh(\frac{\mu}{2})d\bar{v},\quad d\bar{V}=\cosh(\frac{\mu}{2})d\bar{v}+\sqrt{\frac{L_{0}}{\bar{L}_{0}}}\sinh(\frac{\mu}{2})dv,$
$dv=\cosh(\frac{\mu}{2})dV-\sqrt{\frac{\bar{L}_{0}}{L_{0}}}\sinh(\frac{\mu}{2})d\bar{V},\quad d\bar{v}=\cosh(\frac{\mu}{2})d\bar{V}-\sqrt{\frac{L_{0}}{\bar{L}_{0}}}\sinh(\frac{\mu}{2})dV.$
@@
$\bm{\omega}_{\mu}=\left(\cosh(\frac{\mu}{2})-\sinh(\frac{\mu}{2})\right)\bm{\omega}=e^{-\frac{\mu}{2}}\bm{\omega},\quad\bm{\beta}_{\mu}=\cosh(\frac{\mu}{2})\bm{\beta}+\sinh(\frac{\mu}{2})\bm{\bar{{\beta}}}.$ (6.7)
The deformed on-shell action is given by the first term of (2.16):
@@
In the BTZ solution with two complex periods $\bm{\omega}=\omega_{1}+\text{i}\omega_{2},\bm{\beta}=\beta_{1}+\text{i}\beta_{2}$, we have
@@
$\bm{\omega}_{\mu}=\cosh\frac{\mu}{2}\bm{\omega}-\sinh\frac{\mu}{2}\frac{\bm{\beta}\bm{\bar{\omega}}}{\bm{\bar{\beta}}},$ (6.10)
@@
which is related to $I^{[\mu]}_{\text{E,global}}$ simply by $\bm{\tau}_{\mu}\leftrightarrow-\frac{1}{\bm{\tau}_{\mu}}$. Therefore, our on-shell action calculation suggests that the modular transformation of the root-$\overline{\text{T}}$ deformed CFT is the same as the undeformed one.
@@ 7 Discussion
For a 2d CFT, we reproduce the modular transformation (2.17). From the bulk point of view, the transformation rule of the deformation parameter $\mu$ manifests due to its dimension being the same as that of the radial coordinate. We also find that the modular property preserves even when the CFT has a gravitational anomaly and the root-$\overline{\text{T}}$ deformation does not alter the modular transformation. Deriving these properties from the field theory side using the same strategy as the one in [29] would be very interesting. Another closely related possible future direction is to study the modular property of the $\text{J}\bar{\text{T}}$ deformed partition function holographically in the mixed-boundary-condition proposal [112]. The $\text{J}\bar{\text{T}}$ deformed partition function is covariant under modular transformation as shown in [113], in particular the deformation parameter also transforms so we expect that it can also be easily understood from the bulk point of view.
Here we only focus on the classical contribution to the partition function. It is crucial but challenging to include the quantum corrections. For pure AdS gravity theory, the quantum correction arises from summing over states corresponding to the Virasoro descendants. In a recent work [35], this 1-loop effect is studied by assuming that the only effect of the $\overline{\text{T}}$ deformation is to alter the temperature, and then the 1-loop contribution can be computed using the usual heat kernel method on a BTZ geometry. However, as shown in [89], the asymptotic symmetry algebra of the deformed theory is not a naive Virasoro algebra, but actually a state-dependent one. This requires further study from both the field theory and gravity theory perspectives.
In the end, we want to emphasize that the boundary term (2.10) in the deformed on-shell action and the holographic dictionary (1) deserve further exploration. For example, as suggested in [55], it may modify the holographic entanglement entropy. In an oncoming work [114], we also find that naively applying the dictionary to the $\overline{\text{T}}$ deformed cone solution or banana solution [115] will lead to wrong results.
@@ Acknowledgments
We thank Miao He, Yan Liu, Yang Lei and Huajia Wang for the valuable discussion. JT is supported by the National Youth Fund No.12105289 and funds from the University of Chinese Academy of Sciences (UCAS), the Kavli Institute for Theoretical Sciences (KITS). FO would like to thank Masahiro Nozaki very much for his support. The work of FO was supported by funds from UCAS, KITS and The Chinese Academy of Sciences (CAS) President’s International Fellowship Initiative (PI FI).
@@ Appendix A Hubbard-Stratonovich method of $\overline{\text{T}}$
In this appendix, we show some details about the derivation about holographic interpretation of $\overline{\text{T}}$ deformation mainly following [77]. Let $\mathcal{O}_{A}$ be a collection of local operators in CFT which are “single trace” in the sense it is dual to some simple bulk field. For example, the energy-momentum tensor $T_{\alpha\beta}$ is treated as a single trace operator because it is dual to the boundary metric $\gamma_{\alpha\beta}$. Then the double trace deformation can be schematically defined as
@@
Following Hubbard-Stratonovich, we integrate in the fields $h_{A}$ by inserting the resolution of identity
$1=\mathcal{N}\int\mathcal{D}h\exp\left(\frac{N^{2}}{2\mu}\int d^{2}h^{A}(\eta^{-1})_{AB}h^{B}\right)$ (A.3)
@@
Ultimately, we evaluate the $h$ integral with saddle point approximation. Varying the exponent with respect to $h_{A}$ gives
@@
Substituting it into (A.6) we end up with
@@
Note that this is only the result of the first-order deformation. To obtain the deformed theory on the whole $\overline{\text{T}}$ flow we have to work out 1) the deformed source $J^{(\mu)}$; 2) the deformed functional $W^{(\mu)}$. The deformed source $J^{(\mu)}$ has been derived by the variation method we review below. Using the defining property of the generating function $\delta W=\frac{1}{2}\int d^{2}x\sqrt{\gamma}T_{\alpha\beta}\delta\gamma^{\alpha\beta}$, we can obtain a flow equation
@@
The proposal is that $\gamma^{[\mu]}_{\alpha\beta}$ and $T^{[\mu]}_{\alpha\beta}$ are the sources and dual operators of the deformed holographic theory which is still an asymptotic AdS3 gravity theory. The general asymptotic AdS3 solution can be written in the Fefferman-Graham gauge like
@@
To derive the deformed functional, we notice
@@
In the case of $\overline{\text{T}}$ we have
@@
$=$ $W^{(0)}[\gamma_{\alpha\beta}^{[\mu]}=g_{\alpha\beta}^{(0)}+\frac{\mu}{16\pi G_{N}}g_{\alpha\beta}^{(2)}+\frac{\mu^{2}}{(16\pi G_{N})^{2}}g_{\alpha\beta}^{(4)}]-\mu\int d^{2}x\sqrt{\gamma^{(0)}}\text{T$\overline{\text{T}}$ }^{(0)}.$
In the last step we have used the fact $\sqrt{\gamma^{(k)}}\text{T$\overline{\text{T}}$ }^{(k)}$ is a flow invariant and

### v3

@@ 4.2 Thermodynamics
After the $\overline{\text{T}}$ deformation, the entropy of the black hole is not deformed, thus it is equal to101010which is also equal to the black hole entropy of the undeformed theory.
@@
which not only depends on the modular parameter but also depends on the size of the system. For example, the deformation is insignificant when the torus is infinitely long. This is because that the effective deformation parameter is the dimensionless quantity $\rho_{c}/|\bm{\omega}_{\mu}|^{2}$. As expected, the deformed on-shell action still equals to the free energy
@@
even though both sides of this equation are deformed in a non-trivial way. To examine the first law of thermodynamics, we treat $M_{\mu}$, $J_{\mu}$ and $S_{BH}^{[\mu]}$ as functions of $\bm{\beta}_{\mu}$ and $\bm{\bar{\beta}}_{\mu}$ and do the variation with respect to them. It turns out the first law is in the same form of the undeformed one
@@ 4.3 Hawking-Page transition and Modular transformation
By equating them, we find that the Hawking-Page phase transition point remains unchanged at $|\bm{\tau}_{\mu}|=1$. Clearly, the deformed on-shell actions $I_{E,\text{global}}^{[\mu]}$ and $I_{E,\text{BTZ}}^{[\mu]}$ are interrelated through the following modular transformation
@@
This transformation aligns with the modular property exhibited by the $\overline{\text{T}}$ deformed torus partition function, as proved in [29, 30]. To further support our argument regarding modular transformation, let us proceed to compute the deformed on-shell action for other $SL(2,\mathbb{Z})$ black holes. It is noteworthy that since the $\overline{\text{T}}$ deformation alters the temperature, the modular transformation does not commute with the $\overline{\text{T}}$ deformation. In other words, we have
@@
and making the following transformations
@@
we end up with the deformed $SL(2,\mathbb{Z})$ black hole metric which is in the form of (2) with
@@
$\bm{\hat{\omega}}_{\mu}=\bm{\omega}_{\mu}(a_{m}\bm{\tau}_{\mu}+b_{m}),\quad\bm{\hat{\beta}}_{\mu}=\bm{\omega}_{\mu}(c_{m}\bm{\tau}_{\mu}+d_{m}),$ (4.37)
as expected for the $SL(2,{\mathbb{Z}})$ black hole. Since the coordinate transformations (4.35) will not change the on-shell action, we conclude that
@@
which is exactly the modular transformation found in [29]. From the holographic perspective, the modular transformation can be readily understood through the transformations given in equation (4.35). The deformation is irrelevant, and as a result, the deformation parameter $\mu$ has a dimension and scales with the radial coordinate $z$ under a Weyl transformation.
@@
as the modular transformation simply maps a saddle to another one. However, this semi-classical partition function is not holomorphic factorized because $\tau_{\mu}$ and $\bar{\tau}_{\mu}$ transform collectively. To further illustrate that the modular property persists even when the holomorphic and anti-holomorphic sectors are asymmetric, we will explore the $\overline{\text{T}}$ deformation of topologically massive gravity, which is dual to a conformal field theory with distinct holomorphic and anti-holomorphic central charges, i.e. $c_{L}\neq c_{R}$[98, 99]. The AdS/CFT correspondence for TMG is elegantly established in [111]. For the calculation of the deformed on-shell action, we solely require the expression of the holographic energy-stress tensor to define the $\overline{\text{T}}$ deformation, as we shall demonstrate subsequently.
@@ 5 Topologically massive gravity theory
In this section, we consider the Euclidean version of the TMG theory whose action is given by
@@
where $\mathcal{L}_{CS}$ is the gravitational Chern-Simons term. Notably, we assume the parameter $\alpha$ to be real, ensuring that the deformed metric remains real [^12] For the Euclidean TMG, derived through a Wick rotation from the Lorentzian theory, the parameter $\alpha$ is purely imaginary because $\mathcal{L}_{CS}$ has odd parity.. The equation of motion for this theory is as follows
@@
Consequently, the Einstein solutions (2.11) also satisfy the equations of the TMG theory. However, the holographic stress tensor is modified to [99, 106]
@@
where $\epsilon_{ij}$ is the two-dimensional Levi-Civita tensor. Furthermore, the actions of global AdS and BTZ black hole in TMG are given in [112, 113]:
@@
where the involved parameters $k$ and $\bar{k}$ are determined by the relations
@@ 5.1 $\overline{\text{T}}$ deformation
therefore, the flow equation is still solved by (2.4) and (2.5). Substituting the stress tensor (5.5) into (2.4) leads to
$\gamma_{\alpha\beta}^{[\mu]}dx^{\alpha}dx^{\beta}=\left(dv+\frac{\mu\bar{k}}{4\pi}\bar{L}_{0}d\bar{v}\right)\left(d\bar{v}+\frac{\mu k}{4\pi}L_{0}dv\right)\equiv dVd\bar{V},$ (5.9)
@@
Notice that (5.9) can not be interpreted as an induced metric on a cut-off surface, hence, the cut-off (or glue-on) proposal for $\overline{\text{T}}$ deformation is no longer valid. The deformed on-shell action consists of two parts (2.9) which are convenient to express them in terms of the quantities in the seed theory:
@@
where $A_{0}$ is the area of the seed torus. In the end, we will use the dictionary (5.10) to rewrite it in terms of the quantities in the deformed theory. Until now, our discussion is very general. It is enough to demonstrate the modular transformation between the deformed global AdS and deformed BTZ and the other case can be generated trivially as we did in section 4.3.
@@
It has been demonstrated in Section 4.3, the change of the deformation parameter under the modular transformation can be understood from the fact that the $\overline{\text{T}}$ deformation is irrelevant. We proceed to investigate the modular properties of the root-$\overline{\text{T}}$ deformation, which is classically marginal. It is worth noting that the root-$\overline{\text{T}}$ deformation is both compatible with the $\overline{\text{T}}$ deformation and also preserves classical integrability [86]. Although the quantum aspects of root-$\overline{\text{T}}$ deformation are less understood [^14] See [87] for some recent progress., a holographic description has been proposed in [88], which bears similarity to the mixed-boundary-condition description of the $\overline{\text{T}}$ deformation. In particular, the spectrum of the deformed theory is accurately reproduced. The authors of [88] also suggested that the root-$\overline{\text{T}}$ deformed CFT might possess certain modular properties. In particular, in [82] it was shown that S-modular transformation is preserved. We provide some evidences by evaluating the deformed on-shell action, as we did for the $\overline{\text{T}}$ deformation.
@@ 6 Root-$\overline{\text{T}}$ deformation
Because $\mathcal{R}$ is a marginal multi-trace operator, there is no need to add a boundary term to the on-shell action like the one in (2.9). The solution of the flow equation is not unique but it was argued in [88], the proper root-$\overline{\text{T}}$ deformed boundary metric and stress tensor are given by
@@
where we have used the fact that for a classically marginal deformation, the trace of the stress tensor does not flow. Again we can introduce the flat coordinates on the deformed boundary and write the deformed boundary metric as follows
$\gamma_{\alpha\beta}^{(\mu)}dx^{\alpha}dx^{\beta}=dVd\bar{V},$ (6.6)
@@
$dV=\cosh\left(\frac{\mu}{2}\right)dv+\sqrt{\frac{\bar{L}_{0}}{L_{0}}}\sinh\left(\frac{\mu}{2}\right)d\bar{v},\quad d\bar{V}=\cosh\left(\frac{\mu}{2}\right)d\bar{v}+\sqrt{\frac{L_{0}}{\bar{L}_{0}}}\sinh\left(\frac{\mu}{2}\right)dv,$
$dv=\cosh\left(\frac{\mu}{2}\right)dV-\sqrt{\frac{\bar{L}_{0}}{L_{0}}}\sinh\left(\frac{\mu}{2}\right)d\bar{V},\quad d\bar{v}=\cosh\left(\frac{\mu}{2}\right)d\bar{V}-\sqrt{\frac{L_{0}}{\bar{L}_{0}}}\sinh\left(\frac{\mu}{2}\right)dV.$
@@
$\bm{\omega}_{\mu}=\left(\cosh\left(\frac{\mu}{2}\right)-\sinh\left(\frac{\mu}{2}\right)\right)\bm{\omega}=e^{-\frac{\mu}{2}}\bm{\omega},\quad\bm{\beta}_{\mu}=\cosh\left(\frac{\mu}{2}\right)\bm{\beta}+\sinh\left(\frac{\mu}{2}\right)\bm{\bar{{\beta}}}.$ (6.7)
In this case, the deformed on-shell action is given by the first term of (2.16):
@@
On the other hand, in the BTZ solution with two complex periods $\bm{\omega}=\omega_{1}+\text{i}\omega_{2},\bm{\beta}=\beta_{1}+\text{i}\beta_{2}$, we have
@@
${\bm{\omega}_{\mu}=\bm{\omega}\cosh\left(\frac{\mu}{2}\right)-\left(\frac{\bm{\beta}\bm{\bar{\omega}}}{\bm{\bar{\beta}}}\right)\sinh\left(\frac{\mu}{2}\right),}$ (6.10)
@@
which is related to $I^{[\mu]}_{\text{E,global}}$ simply by $\bm{\tau}_{\mu}\leftrightarrow-\frac{1}{\bm{\tau}_{\mu}}$. Notice that in this case, the deformation parameter $\mu$ is not changed under the modular transformation. Therefore, our on-shell action calculation suggests that the modular transformation of the root-$\overline{\text{T}}$ deformed CFT is the same as that of the undeformed one.
@@ 7 Discussion
For a 2d $\overline{\text{T}}$ deformed CFT, we reproduce the modular transformation (2.17) of the partition function. From the bulk point of view, the transformation rule of the deformation parameter $\mu$ arises because its dimension matches that of the radial coordinate. We also find that the modular property (2.17) is preserved even when the CFT has a gravitational anomaly and is dual to a TMG theory. Moreover, we show that the root-$\overline{\text{T}}$ deformation does not alter the behavior of the partition function under the modular transformation, and hence it is the same as that of the undeformed CFT. It should be pointed out that deriving these properties from the field theory side using the same strategy as the one applied in [29] would be very interesting.
Another closely related possible future direction is to study the modular property of the $\text{J}\bar{\text{T}}$ deformed partition function holographically in the mixed-boundary-condition proposal [114]. The $\text{J}\bar{\text{T}}$ deformed partition function is covariant under modular transformation as shown in [115], in particular the deformation parameter also transforms. Therefore, we expect that it can also be easily understood from the bulk point of view.
Here we only focus on the classical contribution to the partition function. It is crucial but challenging to include the quantum corrections. For pure AdS gravity theory, the quantum corrections arise from summing over states corresponding to the Virasoro descendants. In a recent work [36], this 1-loop effect was studied by assuming that the only effect of the $\overline{\text{T}}$ deformation is to alter the temperature, and then the 1-loop contribution can be computed using the usual heat kernel method on a BTZ geometry. However, as shown in [92], the asymptotic symmetry algebra of the deformed theory is not a naive Virasoro algebra, but actually a state-dependent one. Thus, this requires further study from both the field theory and gravity theory perspectives.
In the end, we want to emphasize that the boundary term (2.10) in the deformed on-shell action and the holographic dictionary (1) deserve further exploration. For example, as suggested in [57], it may modify the holographic entanglement entropy. Furthermore, in an oncoming work [116], we also find that naively applying the dictionary to the $\overline{\text{T}}$ deformed cone solution or banana solution [117] will lead to wrong results.
@@ Acknowledgments
We thank Miao He, Yan Liu, Yang Lei and Huajia Wang for the valuable discussion. JT is supported by the National Youth Fund No.12105289 and funds from the University of Chinese Academy of Sciences (UCAS), the Kavli Institute for Theoretical Sciences (KITS). FO would like to thank Masahiro Nozaki very much for his supports. The work of FO was supported by funds from UCAS, KITS and The Chinese Academy of Sciences (CAS) President’s International Fellowship Initiative (PI FI).
@@ Appendix A Hubbard-Stratonovich method of $\overline{\text{T}}$
In this appendix, we show some details about the holographic interpretation of the $\overline{\text{T}}$ deformation mainly following [80]. Let $\mathcal{O}_{A}$ be a collection of local operators in CFT which are “single trace” in the sense it is dual to some simple bulk field. For example, the energy-momentum tensor $T_{\alpha\beta}$ is treated as a single trace operator, because it is dual to the boundary metric $\gamma_{\alpha\beta}$. Then the double trace deformation can be schematically defined as
@@
Following the Hubbard-Stratonovich method, we integrate in the fields $h_{A}$ by inserting the resolution of identity
$1=\mathcal{N}\int\mathcal{D}h\exp\left(\frac{N^{2}}{2\mu}\int d^{2}h^{A}(\eta^{-1})_{AB}h^{B}\right),$ (A.3)
@@
Ultimately, we evaluate the $h$ integral with the saddle point approximation. Varying the exponent with respect to $h_{A}$ gives
@@
Substituting it into (A.6), we end up with
@@
Note that this is only the result of the first-order deformation. To obtain the deformed theory on the whole $\overline{\text{T}}$ flow, we have to work out 1) the deformed source $J^{(\mu)}$; 2) the deformed functional $W^{(\mu)}$. The deformed source $J^{(\mu)}$ has been derived by the variation method that we review below. Using the defining property of the generating function $\delta W=\frac{1}{2}\int d^{2}x\sqrt{\gamma}T_{\alpha\beta}\delta\gamma^{\alpha\beta}$, we can obtain a flow equation
@@
The proposal is that $\gamma^{[\mu]}_{\alpha\beta}$ and $T^{[\mu]}_{\alpha\beta}$ are the sources and dual operators of the deformed holographic theory which is still an asymptotic AdS3 gravity theory. The general asymptotic AdS3 solution can be written in the Fefferman-Graham gauge as follows
@@
To derive the deformed generating functional, we notice
@@
In the case of $\overline{\text{T}}$ deformation, we have
@@
$=$ $W^{(0)}[\gamma_{\alpha\beta}^{[\mu]}=g_{\alpha\beta}^{(0)}+\frac{\mu}{16\pi G_{N}}g_{\alpha\beta}^{(2)}+\frac{\mu^{2}}{(16\pi G_{N})^{2}}g_{\alpha\beta}^{(4)}]-\mu\int d^{2}x\sqrt{\gamma^{(0)}}\text{T$\overline{\text{T}}$ }^{(0)},$
[…]